A particle is oscillating with simple harmonic motion of amplitude 15cm and frequency 4 Hz.

Compute (i) the maximum values of acceleration and velocity, (ii) the acceleration and velocity when

the displacement is 9cm.



1
Expert's answer
2021-11-11T12:00:31-0500

(i) "v_\text{max}=0.15·2\pi·4=3.8\text{ m/s}.\\
a_\text{max}=0.15·(2\pi·4)^2=95\text{ m/s}^2."

(ii) "v=0.09·2\pi·4=2.3\text{ m/s}.\\
a=0.09·(2\pi·4)^2=57\text{ m/s}^2."